# Learning stochastic dynamics from snapshots through regularized unbalanced optimal transport

### v1

@@ Abstract
Reconstructing dynamics using samples from sparsely time-resolved snapshots is an important problem in both natural sciences and machine learning. Here, we introduce a new deep learning approach for solving regularized unbalanced optimal transport (RUOT) and inferring continuous unbalanced stochastic dynamics from observed snapshots. Based on the RUOT form, our method models these dynamics without requiring prior knowledge of growth and death processes or additional information, allowing them to be learnt directly from data. Theoretically, we explore the connections between the RUOT and Schrödinger bridge problem and discuss the key challenges and potential solutions. The effectiveness of our method is demonstrated with a synthetic gene regulatory network. Compared with other methods, our approach accurately identifies growth and transition patterns, eliminates false transitions, and constructs the Waddington developmental landscape.
keywords:
regularized unbalanced optimal transport; Schrödinger bridge; trajectory inference; deep learning.
Contents
@@ 1 Introduction
In machine learning and natural science, a fundamental problem is the coupling of distributions, often high-dimensional, based on the observed samples. For instance, central to generative models as Variational Autoencoders (VAEs) (Kingma and Welling, 2013) is to learn a nonlinear mapping from a complex data distribution to a simpler latent space, where the underlying structure of the data can be more easily captured and generated. In multi-modal data analysis, integration of distribution from disparate data types is essential to provide a more fused and comprehensive understanding of the system (Lahat et al., 2015). Particularly in biology, this has been applied in multi-omics data integration, where distributions from genomics, transcriptomics, proteomics as well as spatial information are aligned into the unified representation of cellular states (Demetci et al., 2022; Cao et al., 2022; Gao et al., 2024b; Cang and Zhao, 2024).
Recently, there has been growing interest in understanding the dynamics of how distributions are coupled over time. In generative machine learning, this is exemplified by diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015), which improves the VAE models by the gradual transformation of a simple distribution into a more complicated data distribution. The diffusion models could be formulated in the framework of stochastic differential equations (SDEs) (Song et al., 2021), and useful for interpolating paths between arbitrary distributions and learning to model the resulting dynamics (Tong et al., 2024b; Neklyudov et al., 2023; De Bortoli et al., 2021; Wang et al., 2021).
The analysis of time series data in single-cell RNA sequencing (scRNA-seq) provides an important application scenario of the high dimensional probability distribution coupling and dynamical inference techniques. Due to the destructive nature of current mainstream sequencing technologies, the continuous trajectories of gene expression dynamics within individual cells are not present in the data. Instead, we can only observe changes in their distributions. Consequently, reconstructing the trajectories of individual cells from samples collected at sparse temporal points in the experiment is crucial for understanding development and other biological processes (Schiebinger et al., 2019; Sha et al., 2024; Peng et al., 2024; Lavenant et al., 2024; Jiang et al., 2022; Jiang and Wan, 2024; Bunne et al., 2024, 2023b; Tong et al., 2023).
@@
The trajectory inference algorithms in scRNA-seq data have been extensively studied (Saelens et al., 2019), and optimal transport (OT)-based methods have emerged as a central tool for datasets with temporal resolution (Bunne et al., 2024). Static OT methods, for instance, have been applied to infer the ancestors or descendants of given cells across time points (Schiebinger et al., 2019; Klein et al., 2023a). Meanwhile, in many cases, there is a need to learn the continuous dynamics of particles over time and fit the mechanistic model that transforms the source distribution into the target distribution. This has led to the development of various dynamical OT methods, as well as flow-based generative methods, including Benamou-Brenier formulation (Benamou and Brenier, 2000) based approaches (Tong et al., 2020; Huguet et al., 2022; Zhang et al., 2024a; Ruthotto et al., 2020; Liu et al., 2021; Cheng et al., 2024a; Wan et al., 2023; Pooladian et al., 2024), unbalanced dynamic OT based approaches (Peng et al., 2024; Sha et al., 2024; Tong et al., 2023), Gromov-Wasserstein (GW) OT approaches (Klein et al., 2023b), continuous normalizing flows (CNF) and conditional flow matching methods (CFM) (Lipman et al., 2023; Tong et al., 2024a; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023b; Jiao et al., 2024b; Gao et al., 2024a; Liu et al., 2022b; Wu et al., 2023; Jin et al., 2024; Chow et al., 2020; Cheng et al., 2024b). Despite the advances, these methods have not fully taken the stochastic dynamical effects into account, especially the intrinsic noise in gene expression and cell differentiation (Zhou et al., 2021a, b), which is prevalent in biological processes on single-cell level (Elowitz et al., 2002).
Pertinent to stochastic dynamics, the Schrödinger bridge (SB) problem aims to identify the most likely stochastic transition path between two arbitrary distributions relative to a reference stochastic process (Léonard, 2014). The SB problem or its variants has been applied in a wide range of context, including scRNA-seq analysis and generative modeling. These approaches, for instance, include static methods (Lavenant et al., 2024; Pariset et al., 2023; Ventre et al., 2023; Chizat et al., 2022; Shi et al., 2024; De Bortoli et al., 2021; Pooladian and Niles-Weed, 2024; Liu et al., 2022a), dynamic methods (Koshizuka and Sato, 2023; Neklyudov et al., 2023, 2024; Zhang et al., 2024b; Bunne et al., 2023a; Chen et al., 2022a; Albergo et al., 2023; Wang et al., 2021; Jiao et al., 2024a; Zhou et al., 2024a; Liu et al., 2023a; Zhou et al., 2024b), and flow-matching methods (Tong et al., 2024b). However, these methods either fail to account for unnormalized distributions resulting from cell growth and death (Sha et al., 2024), or require prior knowledge of these processes (e.g. growth/death rate) (Schiebinger et al., 2019; Lavenant et al., 2024; Pariset et al., 2023; Chizat et al., 2022) or additional information (e.g. cell lineage) (Ventre et al., 2023).
@@
Recently, regularized unbalanced optimal transport (RUOT) offers a promising approach for modeling such stochastic unbalanced continuous dynamics (Baradat and Lavenant, 2021; Chen et al., 2022b; Buze and Duong, 2023; Janati et al., 2020). It can be viewed as an unbalanced relaxation of the dynamic formulation of the Schrödinger bridge problem. In (Baradat and Lavenant, 2021), the authors have provided insights on the connection between certain RUOT and branching Schrödinger bridges. However, computational methods for learning unbalanced stochastic dynamics from samples or such dynamic formulations of certain unbalanced Schrödinger bridges without prior knowledge or additional information are still lacking.
@@
Here we develop a new deep learning method (DeepRUOT) for learning general RUOT and inferring continuous unbalanced stochastic dynamics from samples based on the derived Fisher regularization form without requiring prior knowledge. We demonstrate the effectiveness of our method using a synthetic gene regulatory data. Compared to the common (balanced) Schrödinger bridge method, our RUOT-based approach accurately identifies growth and transition patterns, eliminates false transitions, and constructs the Waddington developmental landscape. On the theoretical side, in this paper we provide a straightforward derivation for the dynamical form of the Schrödinger bridge and then summarize and explore the relationship between RUOT and the Schrödinger bridge problem.
The rest of the paper is organized as follows. In Section 2, we provide preliminaries and background on the Schrödinger bridge and optimal transport, along with a straightforward proof of the dynamical formulation of Schrödinger bridge. Section 3 derives the RUOT problem and its Fisher regularization form. In Section 4, we summarize and explore the relationship between the RUOT problem and the SB problem. Section 5 introduces our deep learning method for learning the RUOT problem from data. In Section 6, we present a synthetic gene regulatory data to validate our framework and demonstrate its effectiveness. We summarize our contributions and discuss possible future directions in Section 7.
@@ 2 Preliminaries and Backgrounds
In this section, we will first review the formulation of the dynamic diffusion Schrödinger Bridge problem. Subsequently, we will provide new proof of the dynamical form of the Schrödinger Bridge. Finally, we will revisit the definitions of the dynamical formulations of optimal transport, regularized optimal transport, and unbalanced optimal transport respectively.
2.1 Schrödinger Bridge Problem
@@
$\min_{\mu^{\boldsymbol{X}}_{0}=\nu_{0},~{}\mu^{\boldsymbol{X}}_{1}=\nu_{1}}\mathcal{D}_{\mathrm{KL}}\left(\mu^{\boldsymbol{X}}_{[0,1]}\|\mu^{\boldsymbol{Y}}_{[0,1]}\right),$ (2.1)
@@
$\mathrm{d}{\boldsymbol{X}}_{t}=\boldsymbol{b}\left({\boldsymbol{X}}_{t},t\right)\mathrm{d}t+\boldsymbol{\sigma}\left({\boldsymbol{X}}_{t},t\right)\mathrm{d}{\boldsymbol{W}}_{t},$ (2.2)
and the reference measure $\mu^{\boldsymbol{Y}}_{[0,1]}$ is chosen as the probability measure induced by the process $\mathrm{d}\boldsymbol{Y}_{t}=\boldsymbol{\sigma}(\boldsymbol{Y}_{t},t)\mathrm{d}{\boldsymbol{W}}_{t}$, where ${\boldsymbol{W}}_{t}\in\mathbb{R}^{d}$ is the standard multidimensional Brownian motion defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ with $\mathbb{P}$ the Wiener measure for the coordinate process ${\boldsymbol{W}}_{t}(\omega)=\omega(t)$, and $\boldsymbol{\sigma}(\boldsymbol{x},t)\in\mathbb{R}^{d\times d}$ is the diffusion matrix which is typically assumed as bounded, coercive and invertible. Under this formulation, the solution to Eq. (2.1) is also referred to as the diffusion Schrödinger bridge. The diffusion Schrödinger bridge also possesses a dynamic formulation, which can be formally stated as the following theorem.
@@
Consider the diffusion Schrödinger bridge problem
$\min_{\mu^{\boldsymbol{X}}_{0}=\nu_{0},~{}\mu^{\boldsymbol{X}}_{1}=\nu_{1}}\mathcal{D}_{\mathrm{KL}}\left(\mu^{\boldsymbol{X}}_{[0,1]}\|\mu^{\boldsymbol{Y}}_{[0,1]}\right),$ (2.3)
where $\mu^{\boldsymbol{Y}}_{[0,1]}$ the reference measure induced by $\mathrm{d}\boldsymbol{Y}_{t}=\boldsymbol{\sigma}(\boldsymbol{Y}_{t},t)\mathrm{d}{\boldsymbol{W}}_{t}$. Then Eq. (2.3) is equivalent to
$\inf_{\left({p},\boldsymbol{b}\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\boldsymbol{b^{T}}(\boldsymbol{x},t)\boldsymbol{a}^{-1}(\boldsymbol{x},t)\boldsymbol{b}(\boldsymbol{x},t)\right]{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (2.4)
@@
$\partial_{t}{p}(\boldsymbol{x},t)=-\nabla_{\boldsymbol{x}}\cdot\left({p}(\boldsymbol{x},t)\boldsymbol{b}(\boldsymbol{x},t)\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\boldsymbol{a}(\boldsymbol{x},t){p}(\boldsymbol{x},t)\right),$ (2.5)
@@
where $\nabla_{\boldsymbol{x}}^{2}:\left(\boldsymbol{a}(\boldsymbol{x},t){p}(\boldsymbol{x},t)\right)=\sum_{ij}\partial_{ij}(a_{ij}p)$ and $\boldsymbol{a}(\boldsymbol{x},t)=\boldsymbol{\sigma}(\boldsymbol{x},t)\boldsymbol{\sigma}^{T}(\boldsymbol{x},t)$, with vanishing boundary condition: $\lim_{|\boldsymbol{x}|\rightarrow\infty}p(\boldsymbol{x},t)=0$.

### v2

@@ Abstract
Reconstructing dynamics using samples from sparsely time-resolved snapshots is an important problem in both natural sciences and machine learning. Here, we introduce a new deep learning approach for solving regularized unbalanced optimal transport (RUOT) and inferring continuous unbalanced stochastic dynamics from observed snapshots. Based on the RUOT form, our method models these dynamics without requiring prior knowledge of growth and death processes or additional information, allowing them to be learned directly from data. Theoretically, we explore the connections between the RUOT and Schrödinger bridge problem and discuss the key challenges and potential solutions. The effectiveness of our method is demonstrated with a synthetic gene regulatory network, high-dimensional Gaussian Mixture Model, and single-cell RNA-seq data from blood development. Compared with other methods, our approach accurately identifies growth and transition patterns, eliminates false transitions, and constructs the Waddington developmental landscape. Our code is available at: https://github.com/zhenyiizhang/DeepRUOT.
@@ 1 Introduction
In machine learning and natural sciences, a key challenge is coupling high-dimensional distributions from observed samples, exemplified by Variational Autoencoders (VAEs) (Kingma & Welling, 2013) which map complex data to simpler latent spaces. It is also important in multi-modal analysis for integrating diverse data types (Lahat et al., 2015), particularly in biology through aligning multi-omics data into unified cellular state representations (Demetci et al., 2022; Cao et al., 2022; Gao et al., 2024b; Cang & Zhao, 2024). Recently, there has been growing interest in understanding the dynamics of how distributions are coupled over time, such as diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015) and stochastic differential equations (SDEs) (Song et al., 2021). The task is useful for interpolating paths between arbitrary distributions and learning the underlying dynamics (Tong et al., 2024b; Neklyudov et al., 2023; De Bortoli et al., 2021; Wang et al., 2021).
@@
Due to the destructive nature of technology, the analysis of time series data in single-cell RNA sequencing (scRNA-seq) provides an important application scenario for the high dimensional probability distribution coupling and dynamical inference problem (Schiebinger et al., 2019; Sha et al., 2024; Peng et al., 2024; Lavenant et al., 2024; Jiang et al., 2022; Jiang & Wan, 2024; Bunne et al., 2024; 2023b; Tong et al., 2023; Zhang et al., 2021). Trajectory inference in scRNA-seq data have been extensively studied (Saelens et al., 2019), and optimal transport (OT)-based methods has emerged as a central tool for datasets with temporal resolution (Bunne et al., 2024; Schiebinger et al., 2019; Klein et al., 2023a; b). Often, there is a need to learn the continuous dynamics of cells over time and fit the mechanistic model that transforms the initial cell distributions into the distributions at later temporal points. This could be solved through the dynamical formulation of OT (Benamou & Brenier, 2000), also known as the B-B form. However, the formulation has not fully taken the stochastic dynamical effects into account, especially the intrinsic noise in gene expression and cell differentiation (Zhou et al., 2021a; b), which is prevalent in biological processes on single-cell level (Elowitz et al., 2002).
@@
By incorporating stochastic dynamics, the Schrödinger bridge (SB) problem aims to identify the most likely stochastic transition path between two arbitrary distributions relative to a reference stochastic process (Léonard, 2014), and has been applied in a wide range of contexts, including scRNA-seq analysis and generative modeling (Pariset et al., 2023; Liu et al., 2022a). Meanwhile, recent regularized unbalanced optimal transport (RUOT) offers a promising approach for modeling both stochastic unbalanced continuous dynamics (Chen et al., 2022b; Baradat & Lavenant, 2021; Buze & Duong, 2023; Janati et al., 2020), which can be viewed as an unbalanced relaxation of the dynamic formulation of Schrödinger bridge problem. However, computational methods for learning RUOT or such high-dimensional unbalanced stochastic dynamics from snapshots are relatively lacking, especially when there is no prior knowledge of unbalanced effect.
@@
Here we develop a new deep learning method (DeepRUOT) for learning general RUOT and inferring continuous unbalanced stochastic dynamics from samples based on the derived Fisher regularization form without requiring prior knowledge. We demonstrate the effectiveness of DeepRUOT on both synthetic and real-world datasets. Compared to the common SB method, our approach accurately identifies growth and transition patterns, eliminates false transitions, and constructs the Waddington developmental landscape of scRNA-seq data. Overall, our main contributions can be summarized as follows:
  • We reformulate RUOT with a Fisher regularization form and explore the connections between RUOT and unbalanced SB. The formulation transforms the SDE into the ordinary differential equation (ODE), which is computationally more tractable.
  • We propose DeepRUOT, the neural network algorithm for learning high dimensional unbalanced stochastic dynamics from snapshots. Through the neural network modeling for growth and death, our framework does not require prior knowledge of these processes.
  • We validate the effectiveness of DeepRUOT on both synthetic data and real scRNA-seq datasets, showing its promising performance compared with existing approaches.
2 Related Works
Deep Learning Solver for Dynamical OT
To tackle the dynamical OT (i.e. B-B form) in high dimensions, many methods (Tong et al., 2020; Huguet et al., 2022; Zhang et al., 2024a; Ruthotto et al., 2020; Liu et al., 2021; Cheng et al., 2024a; Wan et al., 2023; Pooladian et al., 2024; Lipman et al., 2023; Tong et al., 2024a; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023b; Jiao et al., 2024b; Gao et al., 2024a; Liu et al., 2022b; Wu et al., 2023; Jin et al., 2024; Chow et al., 2020; Cheng et al., 2024b) have been developed based on continuous normalizing flow and neural ODE formulation either in original or latent space. To account for sink and source terms in unnormalized distributions, (Sha et al., 2024; Peng et al., 2024; Tong et al., 2023) formulated the neural-network based solver in the unbalanced dynamical OT setup. However, the appropriate formulation along with an effective deep learning solver to simultaneously account for unbalanced term and stochastic effects in dynamical OT remains largely lacking.
Computational Methods for Schrödinger Bridge Problem
Many methods have recently been developed to solve the static SB problem (Lavenant et al., 2024; Pariset et al., 2023; Ventre et al., 2023; Chizat et al., 2022; Shi et al., 2024; De Bortoli et al., 2021; Pooladian & Niles-Weed, 2024; Liu et al., 2022a). To tackle the dynamical Schrödinger Bridge, methods based on neural SDE solver, neural ODE solver with fisher information regularization or flow matching (Koshizuka & Sato, 2023; Neklyudov et al., 2023; 2024; Zhang et al., 2024b; Bunne et al., 2023a; Chen et al., 2022a; Albergo et al., 2023; Wang et al., 2021; Jiao et al., 2024a; Zhou et al., 2024a; Liu et al., 2023a; Zhou et al., 2024b; Tong et al., 2024b; Maddu et al., 2024) have been proposed. However, these methods either fail to account for unnormalized distributions resulting from cell growth and death, or require prior knowledge of the processes (growth/death rate) (Schiebinger et al., 2019; Lavenant et al., 2024; Pariset et al., 2023; Chizat et al., 2022) or additional information (e.g.cell lineage) (Ventre et al., 2023).
Study of RUOT
The RUOT has recently formulated (Chen et al., 2022b), also known as unbalanced Schrödinger Bridge. The existing studies are mostly focused on the theoretical side. For instance, (Janati et al., 2020) derived a closed-form formula for entropic OT between unbalanced Gaussian measures. (Baradat & Lavenant, 2021) investigates relations between RUOT with branching Schrödinger bridge. (Buze & Duong, 2023) investigates different formulations of the problem.
@@ 3 Preliminaries and Backgrounds
In this section, we provide an overview of stochastic effects and unbalanced forms within the dynamical framework. Specifically, considering only stochastic effects leads to the Schrödinger bridge problem (Section 3.1), whereas addressing solely the unbalanced aspect results in unbalanced dynamic optimal transport (Section 3.2). By integrating these two perspectives, we motivate the formulation of the Regularized Unbalanced Optimal Transport (RUOT) framework.
3.1 Stochastic Effect: Schrödinger Bridge Problem
@@
$\min_{\mu^{\boldsymbol{X}}_{0}=\nu_{0},~{}\mu^{\boldsymbol{X}}_{1}=\nu_{1}}\mathcal{D}_{\mathrm{KL}}\left(\mu^{\boldsymbol{X}}_{[0,1]}\|\mu^{\boldsymbol{Y}}_{[0,1]}\right),$ (1)
@@
$\mathrm{d}{\boldsymbol{X}}_{t}=\boldsymbol{b}\left({\boldsymbol{X}}_{t},t\right)\mathrm{d}t+\boldsymbol{\sigma}\left({\boldsymbol{X}}_{t},t\right)\mathrm{d}{\boldsymbol{W}}_{t},$ (2)
and the reference measure $\mu^{\boldsymbol{Y}}_{[0,1]}$ is chosen as the probability measure induced by the process $\mathrm{d}\boldsymbol{Y}_{t}=\boldsymbol{\sigma}(\boldsymbol{Y}_{t},t)\mathrm{d}{\boldsymbol{W}}_{t}$, where ${\boldsymbol{W}}_{t}\in\mathbb{R}^{d}$ is the standard multidimensional Brownian motion defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ with $\mathbb{P}$ the Wiener measure for the coordinate process ${\boldsymbol{W}}_{t}(\omega)=\omega(t)$, and $\boldsymbol{\sigma}(\boldsymbol{x},t)\in\mathbb{R}^{d\times d}$ is the diffusion matrix which is typically assumed as bounded, coercive and invertible. Under this formulation, the solution to Eq. (1) is also referred to as the diffusion Schrödinger bridge. The diffusion Schrödinger bridge also possesses a dynamic formulation, which can be formally stated as the following theorem.
@@
Consider the diffusion Schrödinger bridge problem (1) where $\mu^{\boldsymbol{Y}}_{[0,1]}$ the reference measure induced by $\mathrm{d}\boldsymbol{Y}_{t}=\boldsymbol{\sigma}(\boldsymbol{Y}_{t},t)\mathrm{d}{\boldsymbol{W}}_{t}$. Then (1) is equivalent to
$\inf_{\left({p},\boldsymbol{b}\right)}\int_{0}^{1}\int_{\mathbb{R}^{d}}\left[\frac{1}{2}\boldsymbol{b^{T}}(\boldsymbol{x},t)\boldsymbol{a}^{-1}(\boldsymbol{x},t)\boldsymbol{b}(\boldsymbol{x},t)\right]{p}(\boldsymbol{x},t)\mathrm{d}\boldsymbol{x}\mathrm{d}t,$ (3)
@@
$\partial_{t}{p}(\boldsymbol{x},t)=-\nabla_{\boldsymbol{x}}\cdot\left({p}(\boldsymbol{x},t)\boldsymbol{b}(\boldsymbol{x},t)\right)+\frac{1}{2}\nabla_{\boldsymbol{x}}^{2}:\left(\boldsymbol{a}(\boldsymbol{x},t){p}(\boldsymbol{x},t)\right),$ (4)
@@
where $\nabla_{\boldsymbol{x}}^{2}:\left(\boldsymbol{a}(\boldsymbol{x},t){p}(\boldsymbol{x},t)\right)=\sum_{ij}\partial_{ij}(a_{ij}p)$ and $\boldsymbol{a}(\boldsymbol{x},t)=\boldsymbol{\sigma}(\boldsymbol{x},t)\boldsymbol{\sigma}^{T}(\boldsymbol{x},t)$, coupled with vanishing boundary condition: $\lim_{|\boldsymbol{x}|\rightarrow\infty}p(\boldsymbol{x},t)=0$.